Based on a random sample of fifty full-time college students, we can be 90% confident that for all college students the mean time spent studying per week is between 9.25 hours and 10.75 hours. Which interval (a – d) is a 95% confidence interval for the same sample?

a. (9.10, 10.90)

b. (9.30, 10.70)

c. (9.45, 10.55)

d. (9.00, 10.50)
Answer #1

We know that the 90% confidence interval for population mean is given by:

\\(\overline{x}\pm Z_{0.10/2}\frac{\sigma}{\sqrt{n}})\\ \\ (\overline{x}\pm Z_{0.05}\frac{\sigma}{\sqrt{n}})\\ \\ (\overline{x}\pm (1.645)\frac{\sigma}{\sqrt{n}})\\ \\ \Rightarrow \overline{x}-(1.645)\frac{\sigma}{\sqrt{n}}=9.25, \overline{x}+(1.645)\frac{\sigma}{\sqrt{n}}=10.75

So,

\overline{x}=10,(1.645)\frac{\sigma}{\sqrt{n}}=0.75\Rightarrow \frac{\sigma}{\sqrt{n}}=\frac{150}{329}

We know that the 95% confidence interval for population mean is given by:

\\(\overline{x}\pm Z_{0.05/2}\frac{\sigma}{\sqrt{n}})\\ \\ (\overline{x}\pm Z_{0.025}\frac{\sigma}{\sqrt{n}})\\ \\ (\overline{x}\pm (1.96)\frac{\sigma}{\sqrt{n}})\\ \\ (10\pm (1.96)\frac{150}{329})\\ \\ (10\pm 0.8936)\\ \\(9.10,10.90)

Hence option A.